OpenStudy (anonymous):

There are 5 positions available in the new school. Of the applicants, 12 are men and 8 are women. In how many ways can 3 men and 2 women be chosen if they are equally considered. Select one: a. 6,160 b. 220 c. 28 d. 3,375

OpenStudy (anonymous):

I dont know how to work this one, hopefully @iGreen (the math master) can help you with this one!

OpenStudy (anonymous):

\[\left(\begin{matrix}12 \\ 3\end{matrix}\right)\times \left(\begin{matrix}8 \\ 2\end{matrix}\right) = 6160\]

OpenStudy (anonymous):

there you go^^^^